Week of 180205

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Kate (WLCG, DB, chair), Julia (WLCG), Maarten (Alice, WLCG), Petr (ATLAS), Vincent (security), Alberto (monitoring), Gavin (comp), Borja (mon), Andrei (DB)
  • remote: Xavier (KIT), Andrew (NL-T1), Marcelo (CNAF, LHCb), Dave M (FNAL), Sang Un (KISTI), Di Qing (TRIUMF), Xin Zhao (BNL), Darren (RAL), Christian (NDGF), Christoph (CMS), Kyle (OSG), Pepe (PIC), Vincenzo (EGI)

Experiments round table:

  • ATLAS reports ( raw view) -
    • RAL FTS
      • switched ND, NL and UK clouds to use production instance lcgfts3.gridpp.rl.ac.uk on January 30
      • unfortunately production service run out of available space on Sunday
    • slow transfers to BNL killed after 60 minutes by SRM (< 1MB/s)
      • transfers of big files killed and restarted
      • a lot of rebalancing activity
      • trying to tune site connection limit in FTS (lowered from 700/1000 to 400)
    • EOSATLAS crashed on few times affecting job @ CERN (fixed)
    • CNAF tape library back online - successful transfer from CERN to CNAF tape
Xin commented that BNL is investigating slow transfers. Last week m-core jobs were out. Petr commented there were no inputs for mcore jobs. Parameters on ATLAS side were tuned today and the effects are to be seen. Xin confirmed mcore issue is now gone.

  • CMS reports ( raw view) -
    • No major issues
    • Activity a bit lower compared to previous weeks ~160k cores (~110 Production + ~50k Analysis)
    • Transfers and tape operations remain most demanding items
      • Finishing re-reco of 2017 data
        • Involved staging all all 2017 RAW data and archiving re-reco'd data
      • Continued MC production

  • ALICE -
    • Normal to high activity level on average

  • LHCb reports ( raw view) -
    • Activity
      • HLT farm fully running
      • 2016 data restripping, MC simulation and user jobs
    • Site Issues

Sites / Services round table:

  • ASGC: nc
  • BNL: Slow transfers for ATLAS are being investigated
  • CNAF: Testing access to tape system with personal from Atlas and Alice
    • Wednesday the disks will be tested.
    • More precise news next week
  • EGI: NTR
  • FNAL: NTR
  • IN2P3: NTR
  • JINR: NTR
  • KISTI: NTR
  • KIT: NTR
  • NDGF: NTR
  • NL-T1: NTR
  • NRC-KI: nc
  • OSG: NTR
  • PIC: NTR
  • RAL: There was a problem with our FTS service over the weekend when some disks on the FTS nodes filled up. Atlas submitted a GGUS Alarm ticket yesterday and the problem was resolved by the on-call team.
  • TRIUMF: NTR

  • CERN computing services:NTR
  • CERN storage services: Nothing to report.
  • CERN databases:
    • CMSR database rolling security patch next Wednesday (OTG:0042112)
    • CMSONR, CMSONR ADG and CMSARC database short downtime and a rolling security patch next Wednesday (OTG:0042113)
    • ATLAS ADCR and ATLR databases patched this morning with the latest security updates (OTG:0041967)
    • ATLAS ATONR database will be patched tomorrow morning with the latest security updates (OTG:0042229)
  • GGUS: NTR
  • Monitoring:
    • Not receiving XrootD transfers data from ATLAS US. Fine from other XrootD sources.
  • MW Officer: nc
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r22 < r21 < r20 < r19 < r18 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r22 - 2018-02-05 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback